### v1

@@ Abstract.
Let $G$ be a countable group that admits a minimal equicontinuous action on the Cantor set (i.e., an extension of an infinite residually finite group). We show that the family of almost automorphic actions of $G$ on the Cantor set, is a test for amenability for $G$. More precisely, we show that if $G$ is non-amenable, then for every minimal equicontinuous system $(Y,G)$, such that $Y$ is a Cantor set, there exists a Toeplitz subshift with no invariant probability measures, whose maximal equicontinuous factor is $(Y,G)$. Consequently, we obtain that $G$ is amenable if and only if every Toeplitz $G$-subshift has at least one invariant probability measure.
Key words and phrases:
amenability test, subshifts, almost automorphic, invariant measures.
2020 Mathematics Subject Classification:
Primary: 37B05, 37B10; Secondary: 20E26
@@ 1. Introduction
A classical result states that non-amenable countable groups are characterized as those that admit continuous actions on compact metric spaces with no invariant probability measures (see [1, 17, 33] ). Motivated by this characterization and in response to a question from Grigorchuk, Giordano and de la Harpe show in [18] that to test the non-amenability of a countable group, it suffices to find such an action on the Cantor set. In other words, they show that the class of continuous actions of a countable group $G$ on the Cantor set is a test family for amenability for $G$. A direct application of [18] allows the restriction of the test family even further, for it is sufficient to consider only symbolic systems (see, for example, Lemma 10). Moreover, it is possible to narrow down the size of the alphabet on which the symbolic systems that form the test family for amenability are defined. For example, as a consequence of [27, Theorem 1.2], it can be deduced that for any non-amenable group $G$ there exists a $G$-subshift on $n=n(G)$ letters which admits no invariant probability measure. Recently, Frisch, Kechris, Shinko, and Vidnyánszky show in [14] that the non-amenability of a countable group can be tested within the class of subshifts defined on a two-letter alphabet. Thus, it is natural to wonder whether finding other test families for amenability is possible.
This article focuses on the family of almost automorphic systems, defined as the topological dynamical systems whose maximal equicontinuous factor is minimal and the associated factor map is almost one-to-one (see [15, 29, 30]). The symbolic, almost automorphic systems whose maximal equicontinuous factors are Cantor systems are precisely the Toeplitz $G$-subshifts. These subshifts exhibit a variety of exciting behaviors ([5, 10, 12, 19, 21, 28, 32]), such as the realization of all the metrizable Choquet simplex as spaces of invariant probability measures, and all the possible values for the topological entropy, in the case when the group $G$ is residually finite and amenable (see [7, 11, 24, 32]).
In this article, we show that the family of Toeplitz subshifts is a test for amenability for countable groups which are extensions of infinite residually finite groups. Additionally, this result gives an answer to the question posed by Veech in [30, Page 823] about the existence of almost automorphic systems without invariant measures, which was previously announced by Furstenberg and Weiss in [16].
Our first result is the following:
@@
Let $(Y,G)$ be a minimal topological dynamical system, with $Y$ compact and metric. Let $(Z,G)$ be a minimal equicontinuous system, with $Z$ a Cantor set. Suppose there exists a factor map $\pi:Y\to Z$. Then there exist a minimal topological dynamical system $(X,G)$ and an almost one-to-one factor map $\tau:X\to Z$, such that
  1. (1)
    There exists a factor map $\psi:Y\to X$, such that $\psi$ is one-to-one on $\pi^{-1}(X_{1})$, where $X_{1}\subseteq Z$ is a full measure $G$-invariant set.
@@
    The map $\psi$ induces an affine bijection between the spaces of invariant probability measures of $(Y,G)$ and $(X,G)$.
In particular, $(Y,G)$ has no invariant probability measures if and only if $(X,G)$ has no invariant probability measures.
Furthermore, if $(Y,G)$ is a Cantor minimal system, then so is $(X,G)$.
Theorem 1 was inspired by the techniques developed in [13], [16] and [31, Theorem 5.2]. It can be interpreted as a particular generalization of [13, Theorem 4.3] for group actions.
The following result allows us to get symbolic almost one-to-one extensions with no invariant measures.
Proposition 1.
Let $G$ be a non-amenable countable group. Let $(Y,G)$ be a minimal topological dynamical system, with $Y$ compact and metric, verifying the following properties:
@@
Combining Theorem 1 and Proposition 1, we get the following result:
@@
Let $G$ be a non-amenable group. Then for every minimal equicontinuous system $(Y,G)$, such that $Y$ is a Cantor set, there exists a Toeplitz $G$-subshift with no invariant probability measures, whose maximal equicontinuous factor is $(Y,G)$.
As a corollary, we get the following characterization of amenability:
@@
Let $G$ be a countable group which is an extension of an infinite residually finite group. The following statements are equivalent:
@@
    There exists a minimal equicontinuous system $(Y,G)$, with $Y$ a Cantor set, such that every symbolic almost one-to-one extension of $(Y,G)$ has invariant probability measures.
This article is organized as follows: In Section 2, we recall concepts such as Toeplitz subshifts and odometers and their main associated results. In Section 3, we prove Theorem 1. Finally, in Section 4, we prove Proposition 1, and we deduce Theorem 2 and Corollary 1.
@@ Acknowledgements
The authors warmly thank Raimundo Briceño, Paulina Cecchi, Sebastián Donoso, Tomasz Downarowicz, Patricio Pérez and Samuel Petite, for fruitful discussions.
@@
A topological dynamical system is a pair $(X,G)$ where $X$ is a compact Hausdorff space and $G$ is a countable discrete group acting continuously on $X$. We call $X$ the phase space of the system. For every $x\in X$ and $g\in G$, we denote $gx$ the action of $g$ on $x$. The dynamical system $(X,G)$ is free if $gx=x$ implies $g=1_{G}$, for every $x\in X$. A subset $A\subseteq X$ is $G$-invariant or just invariant, if $gA=A$ for every $g\in G$, where $gA=\{gx:x\in A\}$. The topological dynamical system $(X,G)$ is minimal if for every $x\in X$ its orbit $O_{G}(x)=\{gx:g\in G\}$ is dense in $X$. A minimal component of $(X,G)$ is a closed non-empty $G$-invariant subset $Y$ of $X$ such that $(Y,G)$ is minimal. It is well known that every topological dynamical system has minimal components (see for example [2]). The topological dynamical system $(X,G)$ is equicontinuous if the homeomorphisms defined by the action of $G$, constitutes a uniformly equicontinuous family. When $X$ is a metric space, this is equivalent to say that for every $\varepsilon>0$ there exists $\delta>0$, such that if $d(x,y)<\delta$ then $d(gx,gy)<\varepsilon$, for every $x,y\in X$ and $g\in G$.
Let $(X,G)$ and $(Y,G)$ be two topological dynamical systems. We say that a map $\pi:X\to Y$ is equivariant if $\pi(gx)=g\pi(x)$ for every $g\in G$ and $x\in X$. The equivariant map $\pi$ is a factor map if it is also surjective. If $\pi:X\to Y$ is a factor map, we say that $(X,G)$ is a topological extension or an extension of $(Y,G)$, and $(Y,G)$ is a topological factor or a factor of $(X,G)$. If $\pi$ is bijective, we say that $(X,G)$ and $(Y,G)$ are conjugate. The factor map $\pi:X\to Y$ is almost one-to-one if the set $\{y\in Y:|\pi^{-1}\{\pi(y)\}|=1\}$ is dense in $Y$.
@@ 2.1. Invariant measures
Let $(X,G)$ be a topological dynamical system. We denote by ${\mathcal{M}}(X)$ the space of Borel probability measures of $X$. This space is compact, and metrizable if $X$ also is, with respect to the weak topology (see [9]). An invariant probability measure of $(X,G)$ is a measure $\mu\in{\mathcal{M}}(X)$ such that $\mu(gA)=\mu(A)$, for every $g\in G$ and every Borel set $A\subseteq X$. We denote ${\mathcal{M}}(X,G)$ the space of invariant probability measures of $(X,G)$. It is well known that the group $G$ is amenable if and only if for every topological dynamical system $(X,G)$ the space ${\mathcal{M}}(X,G)$ is non-empty (see, for example, [23]).
The next result is classical in Ergodic Theory (see [9, 26]).
@@
Let $(X,G)$ and $(Y,G)$ be two topological dynamical systems. If $\pi:X\to Y$ is a factor map, then the function $\pi^{*}:{\mathcal{M}}(X)\to{\mathcal{M}}(Y)$ defined as
@@
2.2. Cantor systems and subshifts
We say that the topological dynamical system $(X,G)$ is a Cantor system if $X$ is a Cantor set, i.e., non-empty, compact, metric, totally disconnected and without isolated points. Examples of Cantor systems are given by the full $G$-shifts, which we recall in the next lines.
@@ 2.2. Cantor systems and subshifts
The space $\Sigma^{G}$ is a Cantor set when we endow $\Sigma$ with the discrete topology and $\Sigma^{G}$ with the product topology. The cylinder set given by the finite set $F\subseteq G$ and $\Lambda\in\Sigma^{F}$, is defined as
$[\Lambda,F]=\{x\in\Sigma^{G}:x(g)=\Lambda(g),\mbox{ for every }g\in F\}.$
The topology of $\Sigma^{G}$ is generated by the cylinder sets.
The group $G$ acts continuously on $\Sigma^{G}$ by means of the shift action, which is defined as follows: let $x=(x(g))_{g\in G}\in\Sigma^{G}$ and $h\in G$,
@@
The group $G$ is residually finite if for every $g\in G\setminus\{1_{G}\}$ there exist a finite group $F$ and a group homomorphism $\phi:G\to F$ such that $\phi(g)\neq 1_{F}$. This property is equivalent to the existence of a decreasing sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of finite index subgroups of $G$ such that $\bigcap_{n\in\mathbb{N}}\Gamma_{n}=\{1_{G}\}$. The residual subgroup $R(G)$ of $G$ is the intersection of all the finite index subgroups of $G$. The group $R(G)$ is normal in $G$ and coincides with the intersection of all the finite index normal subgroups of $G$. The group $G$ is residually finite if and only if $R(G)$ is trivial (see [3, Proposition 2.1.11]). Observe that $G/R(G)$ is a residually finite group, for every group $G$. Examples of residually finite groups are the free group $\mathbb{F}_{n}$ with $n$ generators and $\mathbb{Z}^{n}$, for every $n\in\mathbb{N}$.
2.4. Odometers

### v2

@@ Abstract.
Let $G$ be an infinite residually finite group. We show that for every minimal equicontinuous Cantor system $(Z,G)$ with a free orbit, and for every minimal extension $(Y,G)$ of $(Z,G)$, there exist a minimal almost 1-1 extension $(X,G)$ of $(Z,G)$ and a Borel equivariant map $\psi:Y\to X$ that induces an affine bijection $\psi^{*}$ between ${\mathcal{M}}(Y,G)$ and ${\mathcal{M}}(X,G)$, the spaces of invariant probability measures of $(Y,G)$ and $(X,G)$, respectively. If $Y$ is a Cantor set, then $(Y,G)$ and $(X,G)$ are Borel isomorphic, i.e., $\psi^{*}$ is also a homeomorphism. As an application, we show that the family of Toeplitz subshifts is a test for amenability for residually finite groups, i.e., a residually finite group $G$ is amenable if and only if every Toeplitz $G$-subshift has invariant probability measures.
Both authors thanks Lorentz Center, where part of this work was done.
@@ 1. Introduction
Let $G$ be a countable infinite group. In [12] (see also [21]), Furstenberg and Weiss prove that if there is a topological factor map $\pi:Y\to Z$ from a transitive system $(Y,G)$ to a minimal system $(Z,G)$, with $Z$ being totally disconnected, then there exist a minimal dynamical system $(X,G)$, an almost one-to-one factor map $\tau:X\to Z$, and a Borel one-to-one equivariant map $\theta:Y_{0}\subseteq Y\to X$, where $Y_{0}$ is the pre-image by $\pi$ of a full measure set $Z_{0}$ of $Z$, such that $\pi=\tau\circ\theta$. As a consequence, the invariant probability measures of $(X,G)$ supported on $\theta(Y_{0})$ are in a bijection with the invariant probability measures of $(Y,G)$. In other words, the almost 1-1 extension $(X,G)$ has at least as many invariant probability measures as $(Y,G)$. Later, in [10], Downarowicz and Lacroix improve this result for the case $G=\mathbb{Z}$, relaxing some hypotheses on $(Y,\mathbb{Z})$ and $(Z,\mathbb{Z})$, and building a Borel map $\theta$ that induces an affine homeomorphism between the spaces of invariant probability measures of $(Y,\mathbb{Z})$ and $(X,\mathbb{Z})$. In this case, $(Y,\mathbb{Z})$ and the almost 1-1 extension $(X,\mathbb{Z})$ have the same spaces of invariant probability measures.
This article focuses on the case that $G$ is an infinite residually finite group, $(Z,G)$ is a minimal equicontinuous Cantor system with a free orbit, and $(Y,G)$ is a minimal extension of $(Z,G)$. Inspired by the ideas in [12], we show it is possible to construct $(X,G)$ in order that the map $\theta$ induces an affine bijection (a homeomorphism if $Y$ is totally disconnected) between ${\mathcal{M}}(Y,G)$ and ${\mathcal{M}}(X,G)$, the spaces of invariant probability measures of $(Y,G)$ and $(X,G)$, respectively. In particular, if $G$ is non-amenable and $(Y,G)$ has no invariant measures, then neither does $(X,G)$. Thus, as a corollary, we get that if $G$ is non-amenable, then every minimal equicontinuous action of $G$ on the Cantor set with a free orbit has a minimal almost 1-1 extension with no invariant measures. This result answers the question posed by Veech in [20, Page 823] about the existence of almost automorphic systems without invariant measures, which was previously announced by Furstenberg and Weiss in [12], and recently shown by Tsankov [19], using different arguments.
Our first result is stated in the following theorem.
@@
Let $G$ be an infinite residually finite group. Let $(Z,G)$ be a minimal equicontinuous system having a free orbit and such that $Z$ is a Cantor set. Let $(Y,G)$ be a minimal extension of $(Z,G)$ through a topological factor map $\pi:Y\to Z$. Then there exist a minimal extension $(X,G)$ of $(Z,G)$ through an almost 1-1 factor map $\tau:X\to Z$, and a Borel equivariant map $\psi:Y\to X$, satisfying the following:
@@
    $\psi$ is one-to-one on a set $\pi^{-1}(\tilde{Z})$, where $\tilde{Z}\subseteq Z$ is a full-measure set with respect to the unique invariant probability measure in $(Z,G)$.
@@
    The map $\psi$ induces an affine bijection between ${\mathcal{M}}(Y,G)$ and ${\mathcal{M}}(X,G)$, given by $\psi^{*}(\mu)(A)=\mu(\psi^{-1}(A))$, for every $\mu\in{\mathcal{M}}(Y,G)$ and every Borel set $A\subseteq X$.
@@
    If in addition, $Y$ is a Cantor set, then $X$ is also a Cantor set and the map $\psi^{*}$ is an affine homeomorphism between ${\mathcal{M}}(Y,G)$ and ${\mathcal{M}}(X,G)$.
In the context of Theorem 1, points (1) and (2) do not guarantee that $\psi^{*}$ is a bijection between ${\mathcal{M}}(X,G)$ and ${\mathcal{M}}(Y,G)$. Indeed, suppose that $(X,G)$ is a symbolic almost 1-1 extension of $(Z,G)$ having two different ergodic measures $\mu_{1}$ and $\mu_{2}$ with the property that the almost 1-1 factor map $\tau:X\to Z$ is a measure-theoretic conjugacy between $(X,G,\mu_{i})$ and $(Z,G,\nu)$, for $i=1,2$ (such examples exist, see for example, [3] or [22]). Taking $Y=Z$, $\pi=id$ and $\psi=(\tau|_{S_{1}})^{-1}$, with $S_{1}\subseteq X$ a $\mu_{1}$-full measure set such that the restriction of $\tau$ to $S_{1}$ makes a measure-theoretic conjugacy between $(X,G,\mu_{i})$ and $(Z,G,\nu)$, we get conditions (1) and (2) of Theorem 1. Nevertheless, $\psi^{*}$ is not surjective because $\mu_{2}$ (and any other invariant measure $\mu\neq\mu_{1}$) has not pre-image by $\psi^{*}$. Section 4.2 is devoted to show that in our construction, $\psi^{*}$ is a bijection between the spaces of invariant probability measures, and a homeomorphism when the phase spaces are Cantor sets. It is important to note that Theorem 1 is valid for a more restricted class of systems than those in [12] and [21].
As an application of Theorem 1 we get the following result.
@@
Let $G$ be a non-amenable residually finite group. Then for every minimal equicontinuous Cantor system $(Z,G)$ having a free orbit, there exists a Toeplitz $G$-subshift with no invariant probability measures, whose maximal equicontinuous factor is $(Z,G)$.
Theorem 2 shows that the class of Toeplitz subshifts is a test for amenability for residually finite groups. See [11, 13, 18] for more examples of families of dynamical systems which are test for amenability for countable groups.
If $G$ is amenable and residually finite, then every metrizable Choquet simplex is realized as the space of invariant probability measures of a Toeplitz $G$-subshift (see [5, 8]). Theorem 2 can be interpreted as saying that when $G$ is a non-amenable residually finite group, then the empty simplex is realized among the family of Toeplitz $G$-subshifts.
As a corollary of Theorem 2, we get the following characterization of amenability of residually finite groups:
@@
Let $G$ be an infinite residually finite countable group. The following statements are equivalent:
@@
    There exists a minimal equicontinuous Cantor system $(Z,G)$ with a free orbit, such that every symbolic almost one-to-one extension of $(Z,G)$ has invariant probability measures.
This article is organized as follows: in Section 2 we recall basic concepts as odometers, Toeplitz subshifts, equicontinuous systems and residually finite groups. In Section 3 we show that if $(Z,G)$ is a minimal equicontinuous Cantor system with a free orbit, conjugate to the odometer associated to a decreasing sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$ with trivial intersection, then there exists an increasing sequence $(D_{n})_{n\in{\mathbb{N}}}$ of finite subsets of $G$ satisfying several properties, including that every $D_{n}$ is a set of right coset representatives of $G/\Gamma_{n}$. In Section 4 we prove Theorem 1, and in Section 5, we prove Theorem 2 and Corollary 3.
@@ Acknowledgements
The authors warmly thank Sebastián Barbieri, Raimundo Briceño, Paulina Cecchi, Sebastián Donoso, Tomasz Downarowicz, Matthieu Joseph, Patricio Pérez, Samuel Petite, and Todor Tsankov for fruitful discussions.
@@
A topological dynamical system is a pair $(X,G)$ where $X$ is a compact metric space and $G$ is a countable discrete group acting continuously on $X$. We call $X$ the phase space of the system. For every $x\in X$ and $g\in G$, we denote $gx$ the action of $g$ on $x$. The stabilizer of $x\in X$ is the subgroup ${\rm Stab}(x)=\{g\in G:gx=x\}$. The dynamical system $(X,G)$ is free if ${\rm Stab}(x)=\{1_{G}\}$, for every $x\in X$. We say that $(X,G)$ has a free orbit, if there exists $x\in X$ such that ${\rm Stab}(x)=\{1_{G}\}$. A subset $A\subseteq X$ is $G$-invariant or just invariant, if $gA=A$ for every $g\in G$, where $gA=\{gx:x\in A\}$. The topological dynamical system $(X,G)$ is minimal if for every $x\in X$ its orbit $O_{G}(x)=\{gx:g\in G\}$ is dense in $X$. A minimal component of $(X,G)$ is a closed non-empty $G$-invariant subset $Y$ of $X$ such that $(Y,G)$ is minimal. It is well known that every topological dynamical system has minimal components (see for example [1]). The topological dynamical system $(X,G)$ is equicontinuous if for every $\varepsilon>0$ there exists $\delta>0$, such that if $d(x,y)<\delta$ then $d(gx,gy)<\varepsilon$, for every $x,y\in X$ and $g\in G$.
Let $(X,G)$ and $(Y,G)$ be two topological dynamical systems. We say that a map $\pi:X\to Y$ is equivariant if $\pi(gx)=g\pi(x)$ for every $g\in G$ and $x\in X$. The equivariant map $\pi$ is a factor map if it is also continuous and surjective. If $\pi:X\to Y$ is a factor map, we say that $(X,G)$ is a topological extension or an extension of $(Y,G)$, and $(Y,G)$ is a topological factor or a factor of $(X,G)$. If $\pi$ is bijective, we say that $(X,G)$ and $(Y,G)$ are conjugate. The factor map $\pi:X\to Y$ is almost one-to-one if the set $\{y\in Y:|\pi^{-1}\{\pi(y)\}|=1\}$ is dense in $Y$.
@@
Lemma 4.
Let $(Y,G)$ be a minimal topological dynamical system. Let $y_{0}\in Y$. For every $\varepsilon>0$ there exists a finite set $F\subseteq G$, such that for every $y\in Y$ and for every $g\in G$, there exists $h\in gF$ such that $d(h^{-1}y_{0},y)<\varepsilon$.
Proof.
Suppose this is not true. Then there exists $\varepsilon>0$, such that for every finite set $F\subseteq G$, there exist $y\in Y$ and $g\in G$, such that for every $h\in Fg$ we have $d(h^{-1}y_{0},y)\geq\varepsilon$. Let $(F_{n})_{n\in{\mathbb{N}}}$ be an increasing sequence of finite subsets of $G$ such that $1_{G}\in F_{1}$ and whose union is equal to $G$. Let $g_{n}\in G$ and $y_{n}\in Y$ be such that for every $h\in g_{n}F_{n}$ we have $d(h^{-1}y_{0},y_{n})\geq\varepsilon$.
After taking subsequences if necessary, we can assume that $(g_{n}^{-1}y_{0})_{n\in{\mathbb{N}}}$ and $(y_{n})_{n\in{\mathbb{N}}}$ converge to $x\in Y$ and $y\in Y$ respectively. Let $h\in G$. Let $n_{0}\in{\mathbb{N}}$ be such that $h\in F_{n}$, for every $n\geq n_{0}$. We have $g_{n}h\in g_{n}F_{n}$, for every $n\geq n_{0}$. Then
$d(h^{-1}g_{n}^{-1}y_{0},y_{n})\geq\varepsilon,\mbox{ for every }n\geq n_{0}.$
From this we get
$\varepsilon\leq d(h^{-1}g_{n}^{-1}y_{0},y_{n})\leq d(h^{-1}g_{n}^{-1}y_{0},h^{-1}x)+d(h^{-1}x,y)+d(y,y_{n}),$
for every $n\geq n_{0}$, which implies that
$\varepsilon\leq d(h^{-1}x,y)\mbox{ for every }h\in G,$
which is not possible, because $(Y,G)$ is minimal. ∎
@@ 2.1. Invariant measures
Let $(X,G)$ be a topological dynamical system. We denote by ${\mathcal{M}}(X)$ the space of Borel probability measures of $X$. This space is convex, compact, and metrizable with respect to the weak topology (see [7], for details). An invariant probability measure of $(X,G)$ is a measure $\mu\in{\mathcal{M}}(X)$ such that $\mu(gA)=\mu(A)$, for every $g\in G$ and every Borel set $A\subseteq X$. We denote ${\mathcal{M}}(X,G)$ the space of invariant probability measures of $(X,G)$.
It is well known that the group $G$ is amenable if and only if for every topological dynamical system $(X,G)$ the space ${\mathcal{M}}(X,G)$ is non-empty (see, for example, [15]).
The next result is classical in ergodic theory (see, for example, [7]).
@@
Let $(X,G)$ and $(Y,G)$ be topological dynamical systems. If $\pi:X\to Y$ is a factor map, then the function $\pi^{*}:{\mathcal{M}}(X)\to{\mathcal{M}}(Y)$ defined as
@@
2.2. Cantor systems.
We say that the topological dynamical system $(X,G)$ is a Cantor system if $X$ is a Cantor set, i.e., non-empty, compact, metric, totally disconnected and without isolated points. Examples of Cantor systems are given by the $G$-subshifts.
2.2.1. $G$-subshifts.
@@
The space $\Sigma^{G}$ is a Cantor set when we endow $\Sigma$ with the discrete topology and $\Sigma^{G}$ with the product topology. The group $G$ acts continuously on $\Sigma^{G}$ by means of the shift action, which is defined as follows: let $x=(x(g))_{g\in G}\in\Sigma^{G}$ and $h\in G$,
@@
2.2.2. Metric on the space of measures of a Cantor set.
Let $X$ be a Cantor space and let $\mathcal{K}=(\mathcal{K}_{n})_{n\in{\mathbb{N}}}$ be a decreasing sequence of nested clopen partitions of $X$ that spans its topology. Observe that every partition $\mathcal{K}_{n}$ is finite, and the diameter of the elements in $\mathcal{K}_{n}$ goes to zero with $n$. Let define
$\mathcal{S}(\mathcal{K})=\left\{\sum_{K\in\mathcal{K}_{n}}a_{K}\mathbbm{1}_{K}:a_{K}\in{\mathbb{Q}},\mbox{ for every }K\in\mathcal{K}_{n},n\in{\mathbb{N}}\right\}.$
Lemma 7.
For every $\varepsilon>0$ and $f\in C(X)$, there exists $f_{\varepsilon}\in\mathcal{S}(\mathcal{K})$ such that
$\sup_{x\in X}|f(x)-f_{\varepsilon}(x)|<\varepsilon$.
Proof.
Let $\varepsilon>0$, $f\in C(X)$ and $\delta>0$ be such that if $d(x,y)<\delta$ then $|f(x)-f(y)|<\frac{\varepsilon}{2}$. Let $n\in{\mathbb{N}}$ be such that the elements of $\mathcal{K}_{n}$ have diameter smaller than $\delta$. For every $K\in\mathcal{K}_{n}$, we choose $x_{K}\in K$ and $a_{K}\in{\mathbb{Q}}$ such that $|a_{K}-f(x_{K})|<\frac{\varepsilon}{2}$. We define $f_{\varepsilon}=\sum_{K\in\mathcal{K}_{n}}a_{K}\mathbbm{1}_{K}$. Since for every $x\in K$ we have $d(x_{K},x)<\delta$, we get $|f(x)-a_{K}|\leq|f(x)-f(x_{K})|+|f(x_{K})-a_{K}|<\varepsilon$. This implies $\sup_{x\in X}|f(x)-f_{\varepsilon}(x)|<\varepsilon$. ∎
For every $\mu_{1}$ and $\mu_{2}$ in ${\mathcal{M}}(X)$, we define
(2.1) $d(\mu_{1},\mu_{2})=\sum_{n=1}^{\infty}\sum_{K\in\mathcal{K}_{n}}\frac{1}{2^{n}s_{n}}|\mu_{1}(K)-\mu_{2}(K)|,$
where $s_{n}$ is the number of elements in $\mathcal{K}_{n}$, for every $n\in{\mathbb{N}}$. It is straightforward to check that $d$ is a metric on ${\mathcal{M}}(X)$.
Lemma 8.
[…]
Proof.
Let $g\in\mathcal{S}(\mathcal{K})$ and $n\in{\mathbb{N}}$ be such that $g=\sum_{K\in\mathcal{K}_{n}}a_{K}\mathbbm{1}_{K}$, for some $a_{K}\in{\mathbb{Q}}$, for every $K\in\mathcal{K}_{n}$. Let $\delta:=\min\{\frac{\varepsilon}{2^{n}|a_{K}|s_{n}^{2}}\colon K\in\mathcal{K}_{n}\}$. Thus, if $\mu_{1}$ and $\mu_{2}$ are two measures verifying $d(\mu_{1},\mu_{2})<\delta$, then
$\frac{1}{2^{n}s_{n}}|\mu_{1}(K)-\mu_{2}(K)|<d(\mu_{1},\mu_{2})<\frac{\varepsilon}{2^{n}|a_{K}|s_{n}^{2}}\mbox{ for every }K\in\mathcal{K}_{n}.$
This implies
[…]
Corollary 9.
[…]
Proof.
[…]
Corollary 10.
Let $(\mu_{i})_{i\in{\mathbb{N}}}$ be a sequence of measures in ${\mathcal{M}}(X)$ and $\mu\in{\mathcal{M}}(X)$. The following statements are equivalent:
  1. (1)
    The sequence $(\mu_{i})_{i\in{\mathbb{N}}}$ converges weakly to $\mu$;
  2. (2)
    $\lim_{i\to\infty}d(\mu_{i},\mu)=0$.
Proof.
Suppose that $(\mu_{i})_{i\in{\mathbb{N}}}$ converges weakly to $\mu$. Then for every clopen set $C\subseteq X$, we have $\lim_{i\to\infty}\mu_{i}(C)=\mu(C)$. In particular, for every $n\in{\mathbb{N}}$ and $K\in\mathcal{K}_{n}$, we have $\lim_{i\to\infty}\mu_{i}(K)=\mu(K)$. Let $\varepsilon>0$ and $m\in{\mathbb{N}}$ be such that $\sum_{n>m}\frac{1}{2^{n}}<\frac{\varepsilon}{2}.$ Let $i_{0}\in{\mathbb{N}}$ be such that
$|\mu_{i}(K)-\mu(K)|<\frac{\varepsilon}{2}\left(\sum_{n=1}^{m}\frac{1}{2^{n}}\right)^{-1},\mbox{ for every }K\in\mathcal{K}_{j},\mbox{ with }1\leq j\leq m.$
Then, for every $i\geq i_{0}$,
$d(\mu_{i},\mu)$ $=$ $\sum_{n=1}^{\infty}\sum_{K\in\mathcal{K}_{n}}\frac{1}{2^{n}s_{n}}|\mu_{i}(K)-\mu(K)|$
$=$ $\sum_{n=1}^{m}\sum_{K\in\mathcal{K}_{n}}\frac{1}{2^{n}s_{n}}|\mu_{i}(K)-\mu(K)|+\sum_{n>m}\sum_{K\in\mathcal{K}_{n}}\frac{1}{2^{n}s_{n}}|\mu_{i}(K)-\mu(K)|$
$\leq$ $\frac{\varepsilon}{2}+\frac{\varepsilon}{2}$
$=$ $\varepsilon.$
From this we get (1) implies (2).
[…]
This shows $(2)$ implies (1). ∎
@@
The group $G$ is residually finite if for every $g\in G\setminus\{1_{G}\}$ there exist a finite group $F$ and a group homomorphism $\phi:G\to F$ such that $\phi(g)\neq 1_{F}$. This property is equivalent to the existence of a decreasing sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$ such that $\bigcap_{n\in{\mathbb{N}}}\Gamma_{n}=\{1_{G}\}$. Examples of residually finite groups are the free group ${\mathbb{F}}_{n}$ with $n$ generators and $\mathbb{Z}^{n}$, for every $n\in{\mathbb{N}}$ (see [2] for more details about residually finite groups).
2.3.1. Odometers